### v2

@@ 5. Unifying Bayesian, Frequentist, and Imprecise Reasoning
In this section, we show how a diagram involving Bayesian, Conformal, and Imprecise Probabilistic prediction methods, commute in category $\mathbf{UHCont}$. We focus on the latter rather than $\mathbf{WMeas}_{\text{uc}}$ because its objects and morphisms have less restricting requirements.
Since the diagram commuting is an extension of the Full CP Diagram in Theorem 5, we retain assumptions (i)-(iv).
Let $(\Theta,\Sigma_{\Theta})$ be a measurable parameter space, and call $\lambda_{\Theta}$ a fixed $\sigma$-finite Borel measure on $(\Theta,\Sigma_{\Theta})$. Call $\Delta_{\Theta}^{\mathrm{dens}}$ the space of finitely additive probability measures on $\Theta$ admitting a density with respect to $\lambda_{\Theta}$ (Bell and Hagood,, 1988). [^10] This requirement essentially restricts us to focus on countably additive probabilities, see e.g. also here. Endow $\Delta_{\Theta}^{\mathrm{dens}}$ with the $L^{1}(\lambda_{\Theta})$ topology. Consider now the following conditions,
@@
    Measurability. There exists a de Finetti kernel $k_{n}:\Theta\times Y^{n}\rightarrow\mathbb{R}_{+}$ such that, under parameter $\theta$, the joint density of $y^{n}$ is $p(y^{n}\mid\theta)=k_{n}(\theta,y^{n})$ (Hoff,, 2009), and the predictive for $y_{n+1}$ is
    $p_{P}(y_{n+1}\mid y^{n})=\frac{\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))P(\text{d}\theta)}{\int_{\Theta}k_{n}(\theta,y^{n})P(\text{d}\theta)},$
    where $P\in\Delta^{\text{dens}}_{\Theta}$. We assume that $k_{n}(\theta,y^{n})$ is $\Sigma_{\Theta}$-measurable in $\theta$ for each $y^{n}\in Y^{n}$, and Borel in $y^{n}$ for each $\theta$.
@@
    Pointwise $y$-continuity a.e. For $\lambda_{\Theta}$-almost all $\theta$, the map $y^{n}\mapsto k_{n}(\theta,y^{n})$ is continuous on $Y^{n}$.
@@
    Integrable envelope. There exists $M(\theta)\equiv M\in L^{1}(\lambda_{\Theta})$ with $k_{n+1}(\theta,(y^{n},y_{n+1}))\;\leq\;M(\theta)$ for every $(y^{n},y_{n+1})\in Y^{n+1}$ and $\lambda_{\Theta}$-almost all $\theta$.
@@
    Prior integrability. Every admissible prior probability measure $P\in\Delta_{\Theta}^{\mathrm{dens}}$ satisfies $\int_{\Theta}M(\theta)P(\text{d}\theta)<\infty.$ That is, calling $p=\text{d}P/\text{d}\lambda_{\Theta}$ the Radon-Nykodim derivative of $P$ with respect to $\lambda_{\Theta}$, $\int_{\Theta}p(\theta)\,M(\theta)\lambda_{\Theta}(\text{d}\theta)\;<\;\infty$.
  5. (ix)
@@
where $N_{P}(y_{n+1},y^{n})\coloneqq\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta)\lambda_{\Theta}(\text{d}\theta)$. By (ix), it is well-defined.
@@
$\mathsf{BCP}:\;Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\rightarrow Y^{n}\times C(Y^{n}\times Y),\quad\mathsf{BCP}\bigl{(}y^{n},P\bigr{)}\coloneqq\bigl{(}y^{n},\psi_{(y^{n},P)}\bigr{)},$
@@
$\psi_{(y^{n},P)}(y_{n+1})\coloneqq-\,p_{P}(y_{n+1}\mid y^{n});$
We call it $\mathsf{BCP}$, an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma.
@@
Under (v)-(ix), the following hold for every fixed $n\in\mathbb{N}$,
@@
Proof.
Step 1. Continuity of the numerator $N_{P}$. Define
We have $\lvert g\rvert\leq p(\theta)M(\theta)$ by (vii). Assumption (vi) gives continuity of $(y_{n+1},y^{n})\mapsto g(\theta,y_{n+1},y^{n})$ for $\lambda_{\Theta}$-almost all $\theta$. The envelope $pM\in L^{1}(\lambda_{\Theta})$ by (viii); Hence, dominated convergence yields continuity of $(y_{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$.
Step 2. Continuity and positivity of the denominator $D_{P}$. The integrand in $D_{P}(y^{n})$ is $g$, but “stopped” at observation $n$, so the same argument as Step 1 shows that $y^{n}\mapsto D_{P}(y^{n})$ is continuous. Positivity follows from (ix).
$p_{P}(y_{n+1}\mid y^{n})=\frac{N_{P}(y_{n+1},y^{n})}{D_{P}(y^{n})}$
Step 4. Continuity in the prior. Let $P_{m}\to P$ in $L^{1}(\lambda_{\Theta})$, with densities $p_{m}\coloneqq\text{d}P_{m}/\text{d}\lambda_{\Theta}$. From (vii)-(viii)
Consequently
$\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\leq\|p_{P_{m}}(\cdot\mid y^{n}_{m})-p_{P_{m}}(\cdot\mid y^{n})\|_{\infty}+\|p_{P_{m}}(\cdot\mid y^{n})-p_{P}(\cdot\mid y^{n})\|_{\infty}.$
The first term vanishes by uniform continuity from Step 3, the second by Step 4. Hence, $\|\psi_{(y^{n}_{m},P_{m})}-\psi_{(y^{n},P)}\|_{\infty}\to 0$, and the first component $y^{n}_{m}\to y^{n}$ by assumption. Therefore, $\mathsf{BCP}$ is continuous. ∎
@@
Conditions (v)-(ix) are indispensable for Lemma 9 to hold. This is examined in the following remark.
Remark 3 (Minimality of (v)-(ix)).
@@
The five conditions (v)-(ix) are truly minimal for Lemma 9 to hold. If any single hypothesis is dropped, one can construct a Bayesian model for which $\mathsf{BCP}$ fails to be continuous (or even well-defined). To see this, consider the following.
@@
  • Drop (v) (Measurability of $k$). Without measurability in $\theta$, the integrals defining $N_{P}$ and $D_{P}$ need not exist as Lebesgue integrals.
  • Drop (vi) ($y$-continuity a.e.). If $k_{n}(\theta,\cdot)$ has a jump in $y$, then $D_{P}$—and hence $p_{P}$—can jump as a function of $y$, breaking continuity.
@@
  • Drop (viii) (Prior integrability). If $\int_{\Theta}M(\theta)P(\text{d}\theta)=\infty$ for some $P$, then $N_{P}$ or $D_{P}$ may diverge or lose continuity.
  • Drop (ix) (Positive marginal likelihood). If $D_{P}(y^{n})=0$ for some $y^{n}$, the posterior density $p_{P}$ is undefined there, so $\mathsf{BCP}$ cannot be continuous.
In practice, conditions (v)-(ix) are almost built-in to any reasonable Bayesian model. Let us briefly expand on this. First, notice that almost every statistical model (Gaussian, exponential family, mixture, etc.) has a likelihood density continuous and measurable in $y$, so that (v) is met.
Condition (vi) guarantees that, as we vary our data $y^{n}$, the integrand in $N_{P}$ and $D_{P}$ moves continuously, so dominated convergence applies. In practice, standard density kernels (Normal, Laplace, Student-$t$, etc.) are smooth in $y$.
The uniform-in-$y$ bound of (vii) is the key to swapping limits and integrals. If $Y$ is compact (as it is in our setting, by (i)) and $k_{n}(\theta,\cdot)$ is continuous, we can often take $M(\theta)=\sup_{y\in Y}k_{n}(\theta,y)$, which is finite and $L^{1}$-integrable under mild assumptions on $k_{n}$.
Since $N_{P}$ involves $k_{n+1}$, the moment condition in (viii) ensures finiteness and continuity of the predictive density. Common priors (Gaussian, Gamma, Beta, compact support) have finite moments of all orders, so (viii) is automatic.
Condition (ix) prevents division by zero in the posterior $p_{P}(y_{n+1}\mid y^{n})=N_{P}/D_{P}$. In well-posed Bayesian models, the prior’s support overlaps every parameter value that could generate the observed data, so $D_{P}$ never vanishes.
It is safe to say, then, that hypotheses (v)-(ix) are the minimal technical machinery needed to ensure the posterior predictive is well-defined and varies continuously in both the data $y^{n}$ and the prior $P$. In practical modeling, they hold almost automatically.
Before the main result of this section, we need to introduce a further condition on the posterior predictive Cumulative Density Function (CDF).
  1. (x)
    Well-behaved CDF. For each fixed $(y^{n},P)\in Y^{n}\times\Delta_{\Theta}^{\text{dens}}$, the CDF
    $F(c;\,y^{n},P)\coloneqq\int_{\{\,p_{P}(\cdot\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\text{d}y$
    is continuous and strictly increasing in a neighborhood of its $(1-\alpha)$-quantile
    $c(\alpha,y^{n},P)=\inf\{\,c:F(c;y^{n},P)\geq 1-\alpha\}.$
@@
Under assumptions (i)-(x), the correspondence $\mathsf{QUANT}$ is upper hemicontinuous.
Proof.
First, we show that under (i)-(x), the correspondence $H$ is upper hemicontinuous. Fix $(c_{0},y^{n}_{0},P_{0})$ and let $V\supset H(c_{0},y^{n}_{0},P_{0})$ be open in $Y$. Since $p_{P}(y\mid y^{n})$ is jointly continuous and $Y$ is compact, $H(c_{0},y^{n}_{0},P_{0})=\{y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}\}$ is closed and contained in $V$. Thus there is $\varepsilon>0$ such that
$\bigl{\{}y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}-\varepsilon\bigr{\}}\;\subset\;V.$
By joint continuity of $(c,y^{n},P,y)\mapsto p_{P}(y\mid y^{n})$, there is a neighborhood $U$ of $(c_{0},y^{n}_{0},P_{0})$ in ×$\mathbb{R}\times Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}$ such that for all $(c,y^{n},P)\in U$ and all $y\in Y$,
$p_{P}(y\mid y^{n})\;\geq\;c\quad\Longrightarrow\quad p_{P_{0}}(y\mid y^{n}_{0})\;\geq\;c_{0}-\varepsilon.$
Hence for every $(c,y^{n},P)\in U$,
$H(c,y^{n},P)=\{y:p_{P}(y\mid y^{n})\geq c\}\;\subset\;\{y:p_{P_{0}}(y\mid y^{n}_{0})\geq c_{0}-\varepsilon\}\;\subset\;V,$
proving upper hemicontinuity.
Now, we show that under (i)-(x), function $c$ is continuous. By (x), for each fixed $(y^{n},P)$ the function $c\mapsto F(c;y^{n},P)$ is continuous and strictly increasing near the target quantile. Standard results on the continuity of inverse functions then imply that
$(\alpha,y^{n},P)\;\longmapsto\;c(\alpha,y^{n},P)$
is continuous at every point, since it solves $F\bigl{(}c(\alpha,y^{n},P);y^{n},P\bigr{)}=1-\alpha$ uniquely in $c$, and both the left side and the right side vary continuously in $(\alpha,y^{n},P)$.
@@
As we showed before, $H$ is upper hemicontinuous in its three arguments, and $c(\alpha,y^{n},P)$ is continuous, hence the singleton-valued correspondence associated with $c$ (that we still denote by $c$ for ease of notation) is upper hemicontinuous. The composition of two upper hemicontinuous correspondences is again upper hemicontinuous (Aliprantis and Border,, 2006, Theorem 17.23). Therefore, $\mathsf{QUANT}$ is upper hemicontinuous. ∎
Let us add a brief discussion on condition (x).
Remark 4 (On the Well-Behavior of the Predictive CDF).
In many continuous Bayesian models, $p_{P}(y\mid y^{n})$ is a continuous, strictly positive function on compact $Y$ (for us, this holds by (i)). As $c$ varies, the level-sets $\{y:p_{P}(y\mid y^{n})\leq c\}$ move continuously, so $F(c)$ has no jumps. Since $p_{P}$ has no plateaus of positive volume, for any $c<c^{\prime}$, the set $\{c<p_{P}(y\mid y^{n})\leq c^{\prime}\}$ has strictly positive mass, forcing $F(c^{\prime})-F(c)>0$. Hence, $F(c)$ is both continuous and strictly increasing at the target quantile.
The condition may fail in two cases, rather negligible in practice. If $p_{P}(y\mid y^{n})\equiv c$ on a subset of $Y$ of positive measure, then $F(c)$ is constant across that plateau, violating strict increase. Also, if the predictive has atoms (point-mass components), $F(c)$ jumps at those mass points and fails to be continuous.
Lemmas 9 and 10 are crucial in light of the following, which is the main result of this section.
@@
Assume that (i)-(x) hold. Consider the restrictions $\kappa_{\alpha}$, $\mathsf{IHDR}_{\alpha}$, and $\mathsf{QUANT}_{\alpha}$ to any $\alpha\in[0,1]$ that satisfies (iii). Then, the following diagram
@@
commutes in $\mathbf{UHCont}$, that is, $\mathsf{QUANT}_{\alpha}=\kappa_{\alpha}\circ\mathsf{BCP}=\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}\circ\mathsf{BCP}$.
Proof.
First, notice that we do not require consonance directly because it is immediately satisfied as a result of (i) and $(y_{n+1},y^{n})\mapsto p_{P}(y_{n+1}\mid y^{n})$ being jointly continuous on $Y^{n}\times Y=Y^{n+1}$. They imply the existence of a maximizer $y^{\star}\in Y$ such that
$p_{P}(y^{\star}\mid y^{n})=\max_{y\in Y}p_{P}(y\mid y^{n}).$
Hence, $\psi_{(y^{n},P)}(y_{i})\geq\psi_{(y^{n},P)}(y^{\star})$, for every $i\in\{1,\dots,n\}$, and likewise for the candidate $y_{n+1}=y^{\star}$. Therefore every one of the $n\!+\!1$ indicator terms inside the transducer equals $1$, so
$\pi_{\psi_{(y^{n},P)}}(y^{\star},y^{n})=\frac{1}{n+1}\sum_{i=1}^{n+1}1=1.$
Consequently, $\sup_{y\in Y}\pi_{\psi_{(y^{n},P)}}(y,y^{n})=1$, that is, consonance holds automatically under (i) and the (joint) continuity of $p_{P}(y\mid y^{n})$.
Notice also that, by (i) and $\Delta^{\text{dens}}_{\Theta}$ being endowed the $L^{1}(\lambda_{\Theta})$ topology, $Y^{n}\times\Delta^{\text{dens}}_{\Theta}$ is a topological space in the product topology, hence a well-defined object of $\mathbf{UHCont}$.
In addition, by Lemma 9, function $\mathsf{BCP}$ is continuous. The latter implies that the singleton-valued correspondence associated with $\mathsf{BCP}$ (that we still denote by $\mathsf{BCP}$ for ease of notation) is upper hemicontinuous. Moreover, by Lemma 10, correspondence $\mathsf{QUANT}$ (and also, trivially, its restriction $\mathsf{QUANT}_{\alpha}$ to some $\alpha$ satisfying (iii)) is upper hemicontinuous as well. In turn, they are well-defined morphisms of $\mathbf{UHCont}$.
Furthermore, we have that $t(\mathsf{BCP})=s(\kappa_{\alpha})=s(\mathsf{CRED})=Y^{n}\times\mathscr{F}$. That is, the target space of $\mathsf{BCP}$ coincides with the source spaces of both $\kappa_{\alpha}$ and $\mathsf{CRED}$. This is a consequence of

### v3

@@ 4. Unifying Bayesian, Frequentist, and Imprecise Reasoning
In this section, we show how a diagram involving Bayesian, Conformal, and Imprecise Probabilistic prediction methods, commutes asymptotically in category $\mathbf{UHCont}$. Since the diagram commuting is an extension of the Full CP Diagram in Theorem 5, we retain assumptions (i)-(iv).
Let $(\Theta,\Sigma_{\Theta})$ be a measurable parameter space, and call $\lambda_{\Theta}$ a fixed $\sigma$-finite Borel measure on $(\Theta,\Sigma_{\Theta})$. Call $\Delta_{\Theta}^{\mathrm{dens}}$ the space of finitely additive probability measures on $\Theta$ admitting a density with respect to $\lambda_{\Theta}$ (Bell and Hagood,, 1988). [^10] This requirement essentially restricts us to focus on countably additive probabilities, see e.g. also here. Endow $\Delta_{\Theta}^{\mathrm{dens}}$ with the $L^{1}(\lambda_{\Theta})$ topology. Denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$. Consider now the following conditions,
@@
    Measurability. There exist de Finetti kernels $k_{\iota}:\Theta\times Y^{\iota}\to\mathbb{R}_{+}$, $\iota\in\{n,n+1\}$, such that, under parameter $\theta$, the joint density of $y^{\iota}$ is $k_{\iota}(\theta,y^{\iota})$ (a density on $Y^{\iota}$ with respect to $\lambda_{Y}^{\otimes\iota}$). For each $\iota\in\{n,n+1\}$ and each $y^{\iota}\in Y^{\iota}$, $\theta\mapsto k_{\iota}(\theta,y^{\iota})$ is $\Sigma_{\Theta}$-measurable, and for each $\theta\in\Theta$, $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ is Borel. The posterior predictive is
    $p_{P}(y_{n+1}\mid y^{n})=\frac{\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))P(\mathrm{d}\theta)}{\int_{\Theta}k_{n}(\theta,y^{n})P(\mathrm{d}\theta)},\quad P\in\Delta_{\Theta}^{\mathrm{dens}}.$
@@
    Pointwise $y$-continuity a.e. For $\lambda_{\Theta}$-almost all $\theta$, the maps $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ are continuous on $Y^{\iota}$, for $\iota\in\{n,n+1\}$.
@@
    Integrable envelope. There exists $M(\theta)\equiv M\in L^{\infty}(\lambda_{\Theta})$ with $k_{\iota}(\theta,y^{\iota})\leq M(\theta)$ for every $y^{\iota}\in Y^{\iota}$, $\iota\in\{n,n+1\}$, and $\lambda_{\Theta}$-almost all $\theta$.
@@
where $N_{P}(y_{n+1},y^{n})\coloneqq\int_{\Theta}k_{n+1}(\theta,(y^{n},y_{n+1}))p(\theta)\lambda_{\Theta}(\text{d}\theta)$. By (viii), it is well-defined.
@@
$\mathsf{BCP}:Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}\rightarrow Y^{n}\times C(Y^{n}\times Y),\quad\mathsf{BCP}\bigl{(}y^{n},P\bigr{)}\coloneqq\bigl{(}y^{n},\psi_{(y^{n},P)}\bigr{)},$ (10)
@@
$\psi_{(y^{n},P)}(y_{n+1})\coloneqq-p_{P}(y_{n+1}\mid y^{n}).$
Here we abuse notation, and write $\psi_{(y^{n},P)}(y_{n+1})\equiv\psi_{P}(y^{n},y_{n+1})$ to highlight the dependence of the non-conformity score on the collected dataset $y^{n}$. We call $\mathsf{BCP}$ the function in (10), an acronym for Bayesian Conformal Prediction (Fong and Holmes,, 2021). Next, we have an important lemma, that we prove in Appendix B.
@@
Under (i), (v)-(viii), the following hold for every fixed $n\in\mathbb{N}$,
@@
Conditions (v)-(viii) are indispensable for Lemma 7 to hold. This is examined in the following remark.
Remark 2 (Minimality of (v)-(viii)).
@@
The four conditions (v)(viii) are minimal for Lemma 7. If any single hypothesis is dropped, we can construct a Bayesian model for which $\mathsf{BCP}$ fails to be continuous (or even well-defined). In particular,
@@
  • Drop (v) (measurability of $k_{\iota}$). Without $\Sigma_{\Theta}$-measurability in $\theta$, the integrals defining $N_{P}$ and $D_{P}$ may not be Lebesgue integrals. Without Borel measurability in $y^{\iota}$, $(y^{n+1},y^{n})\mapsto N_{P}(y_{n+1},y^{n})$ and $y^{n}\mapsto D_{P}(y^{n})$ need not even be measurable, so continuity cannot hold.
  • Drop (vi) (pointwise $y$-continuity a.e.). Steps 1-2 in the proof require continuity in the data arguments: For $\lambda_{\Theta}$-almost all $\theta$, $y^{\iota}\mapsto k_{\iota}(\theta,y^{\iota})$ must be continuous. If $k_{n+1}(\theta,\cdot)$ has a jump, then $N_{P}$ can jump in $(y_{n+1},y^{n})$; If $k_{n}(\theta,\cdot)$ has a jump, then $D_{P}$ can jump in $y^{n}$. Either breaks continuity of $p_{P}=N_{P}/D_{P}$.
  • Drop (vii) (integrable envelope). The dominated convergence steps (for $N_{P}$ and $D_{P}$ as functions of the data) and the uniform $L^{1}$ control in the prior argument rely on an integrable envelope $M\in L^{\infty}(\lambda_{\Theta})$ bounding $k_{n+1}$ and $k_{n}$, for almost all $\theta$. Without such $M$, we can have $P_{m}\to P$ in $L^{1}$ but $N_{P_{m}}\not\to N_{P}$ (and similarly for $D_{P}$), so $\mathsf{BCP}$ is discontinuous.
  • Drop (viii) (positive marginal likelihood). If $D_{P}(y^{n})=0$ for some $y^{n}$, then $p_{P}(\cdot\mid y^{n})$ is undefined (division by zero), so $\mathsf{BCP}$ is not even well-defined there, let alone continuous.
In practice, conditions (v)(viii) are almost built in to reasonable Bayesian models. First, in standard dominated parametric families (Gaussian, exponential families, finite mixtures with continuous component densities), the likelihood kernel $f_{\theta}(y)$ is continuous in $y$, and the joint density $k_{n}(\theta,y^{n})=\prod_{i=1}^{n}f_{\theta}(y_{i})$ is Borel in $y^{n}$ and $\Sigma_{\Theta}$-measurable in $\theta$, so (v) holds.
Condition (vi) ensures that, as the data $y^{n}$ vary, the integrands in $N_{P}$ and $D_{P}$ vary continuously (for $\lambda_{\Theta}$-almost all $\theta$), so dominated convergence applies. In practice, standard kernels (Normal, Laplace, Student-$t$, etc.) are smooth in $y$.
The uniform-in-$y$ bound in (vii) is what allows swapping limits and integrals in the proof of Lemma 7. Because $Y$ is compact by (i) and $k_{\iota}(\theta,\cdot)$ is continuous, one can take
$M_{n}(\theta)\coloneqq\sup_{y^{n}\in Y^{n}}k_{n}(\theta,y^{n}),\qquad M_{n+1}(\theta)\coloneqq\sup_{y^{n+1}\in Y^{n+1}}k_{n+1}(\theta,y^{n+1}),$
and use $M(\theta)\geq\max\{M_{n}(\theta),M_{n+1}(\theta)\}$. Under mild conditions on $\Theta$ (so that the kernels are uniformly bounded), $M\in L^{\infty}(\lambda_{\Theta})$, so (vii) holds. [^11] Alternatively, we can ask that $M$ is in $L^{1}$, and topologize $\Delta^{\text{dens}}_{\Theta}$ with the weighted topology $L^{1}(M\text{d}\lambda_{\Theta})$.
Since $N_{P}$ involves $k_{n+1}$, the prior integrability condition $\int M(\theta)P(\mathrm{d}\theta)<\infty$ guarantees finiteness and continuity of the predictive. This is usually automatic with common priors when the envelope $M$ has at most moderate growth in $\theta$ (e.g., exponential-family models with conjugate priors).
Finally, (viii) rules out division by zero in $p_{P}(y_{n+1}\mid y^{n})=N_{P}/D_{P}$. In well-posed models, the prior assigns positive mass to parameter values that could have generated the observed data (i.e., $k_{n}(\theta,y^{n})>0$ on a set of $\theta$ with positive prior mass), ensuring $D_{P}(y^{n})>0$.
In short, (v)(viii) are minimal technical machinery to ensure the posterior predictive is well defined and depends continuously on both the data $y^{n}$ and the prior $P$, and they are typically satisfied in practical modeling. $\triangle$
Before the main result of this section, we need to introduce two further conditions, one of which on the posterior predictive Cumulative Density Function (CDF). Recall that we denote by $\Delta_{Y}^{\text{dens}}$ the space of finitely additive probability measures on $Y$ that admit a density with respect to a $\sigma$-finite dominating (atomless) measure $\lambda_{Y}$.
  1. (ix)
    Well-behaved CDF (uniform near the target). Fix $\alpha\in(0,1)$. For each $(y^{n},P)\in Y^{n}\times\Delta_{\Theta}^{\mathrm{dens}}$, define
    $F(c;y^{n},P)\coloneqq\int_{\{y:p_{P}(y\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\lambda_{Y}(\mathrm{d}y),\quad c(\alpha,y^{n},P)\coloneqq\inf\{c:F(c;y^{n},P)\geq\alpha\}.$
@@
    $\big{[}c(\alpha,y^{n},P)-\delta,c(\alpha,y^{n},P)+\delta\big{]}.$
[…]
@@
Under assumptions (i)-(ix), the correspondence $\mathsf{QUANT}$ is upper hemicontinuous.
Lemma 8 is proved in Appendix B. Let us add a brief discussion on condition (ix). In many continuous Bayesian models, the posterior predictive density $y\mapsto p_{P}(y\mid y^{n})$ is continuous and strictly positive on compact $Y$. Then, for every $c$ with $\lambda_{Y}\left(\{y:p_{P}(y\mid y^{n})=c\}\right)=0$, the map
$F(c;y^{n},P)=\int_{\{p_{P}(\cdot\mid y^{n})\leq c\}}p_{P}(y\mid y^{n})\lambda_{Y}(\mathrm{d}y)$
is continuous (by dominated convergence). Moreover, $F$ is strictly increasing at a level $c_{0}$ whenever $c_{0}$ lies in the interior of the essential range of $p_{P}(\cdot\mid y^{n})$, i.e., there exists $\varepsilon>0$ such that
$\lambda_{Y}\left(\{c_{0}-\varepsilon<p_{P}(y\mid y^{n})\leq c_{0}\}\right)>0\quad\text{and}\quad\lambda_{Y}\left(\{c_{0}<p_{P}(y\mid y^{n})\leq c_{0}+\varepsilon\}\right)>0.$
In particular, if $c_{0}=c(\alpha,y^{n},P)$ satisfies this condition, then $F(\cdot;y^{n},P)$ is continuous and strictly increasing on $c_{0}-\varepsilon,c_{0}+\varepsilon]$, so (ix) holds. This is automatic, for example, if $Y$ is connected and $\lambda_{Y}$ assigns positive measure to every nonempty open set.
Condition (ix) may fail if: (i) $p_{P}$ has a plateau of positive $\lambda_{Y}$-measure at some level $c_{\star}$ (so $F$ is locally flat there); (ii) the predictive is not absolutely continuous with respect to $\lambda_{Y}$ (atoms create jumps of $F$); or (iii) $p_{P}$ never attains values in a neighborhood of $c(\alpha,y^{n},P)$ (a “gap” in its range, e.g., on disconnected $Y$), which precludes local strict increase.
The last condition that we need is the following.
  1. (x)
    Leave-One-Out (LOO) stability and conditional Glivenko-Cantelli (GC). Recall that $\mathfrak{P}$ denotes the sampling law of the exchangeable process $(\mathcal{Y}_{i})_{i\geq 1}$. For $(y^{n},P)$ and $y\in Y$, define
    $r_{n}(y)\coloneqq p_{P}(y\mid y^{n}),\qquad r_{n,-i}(y)\coloneqq p_{P}(y\mid y^{n+1}_{-i}),\quad 1\leq i\leq n.$
    1. (a)
      LOO stability.
      $\mathsf{err}_{n}(y^{n},P)\coloneqq\sup_{y\in Y}\max_{1\leq i\leq n}\bigl{|}r_{n,-i}(y)-r_{n}(y)\bigr{|}=o_{\mathfrak{P}}(1).$
    2. (b)
      Conditional GC. Let
@@
      We have
      $\sup_{t\in\mathbb{R}}\bigl{|}F_{n}(t;y^{n},P)-F(t;y^{n},P)\bigr{|}=o_{\mathfrak{P}}(1).$
A natural question at this point is why we need to introduce condition (x). Full CP compares $r_{n,-i}(y_{i})$ with $r_{n}(y)$; The left-hand side depends on the candidate $y$ through $y^{n+1}_{-i}$. To recover the population $\alpha$-level set threshold defined by $F(\cdot;y^{n},P)$, we need two probabilistic approximations, $\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\{r_{n,-i}(y_{i})\leq r_{n}(y)\}]\approx\frac{1}{n}\sum_{i=1}^{n}\mathbbm{1}[\{r_{n}(y_{i})\leq r_{n}(y)\}]\approx F\bigl{(}r_{n}(y);y^{n},P\bigr{)},$ uniformly in $y$. The first approximation is the LOO stability (x).(a); The second is the conditional GC property (x).(b). Neither follows from (i)(ix) (which are analytic/topological).
Conditions like (x) are mild and hold in common settings: (i) In regular parametric Bayes (e.g., smooth exponential families on compact $Y$ with conjugate priors), $r_{n,-i}$ and $r_{n}$ differ by $O_{\mathfrak{P}}(1/n)$ uniformly in $y$ (posterior contraction / Lipschitz dependence on empirical statistics), yielding (x).(a); (ii) In regular dominated models on compact $Y$ (parametric or suitably controlled nonparametric), the posterior predictives $r_{n}$ converge uniformly to a continuous limit $r$, and the posterior predictive laws $Q_{n}(\cdot)\coloneqq\int_{(\cdot)}p_{P}(y\mid y^{n})\lambda_{Y}(\text{d}y)$ converge in probability (as random probability measures) to a probability measure $Q$ with density $r$, i.e. $\|r_{n}-r\|_{\infty}\xrightarrow{\mathfrak{P}}0$, $Q_{n}\to Q\text{ in probability}$, $\frac{\text{d}Q}{\text{d}\lambda_{Y}}=r$. If, in addition, there are no plateaus at the target level (equivalently, $Q(\{r=t\})=0$ for the relevant thresholds $t$), then (x).(b) holds by a uniform Law of Large Numbers applied to the sublevel sets $\{r\leq t\}$. Strengthening both to almost sure versions upgrades all conclusions from “in probability” to “almost surely”.
Lemmas 7 and 8 are crucial in light of the following, which is the main result of this section.
@@
Assume that (i)-(x) and consonance hold. Fix any $\alpha\in(0,1)$ and define $\beta_{n}(\alpha)\coloneqq\min\{k/(n+1):k/(n+1)>\alpha\}$. Then,
$d_{H}\left(\kappa_{\alpha}(y^{n},\psi_{(y^{n},P)}),\mathsf{QUANT}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Consequently, since $\mathsf{IHDR}_{\alpha}\circ\mathsf{CRED}=\kappa_{\alpha}$ under consonance,
$d_{H}\left(\mathsf{IHDR}_{\alpha}\bigl{(}\mathsf{CRED}(y^{n},\psi_{(y^{n},P)})\bigr{)},\mathsf{QUANT}_{\alpha}(y^{n},P)\right)\xrightarrow{\mathfrak{P}}0.$
Equivalently, the diagram
@@
commutes asymptotically in probability in $\mathbf{UHCont}$ (in the Hausdorff distance $d_{H}$).
If $Y$ is compact metric, [^12] We also tacitly assume that the level set $\{y:r_{n}(y)=c(\alpha,y^{n},P)\}$ has empty interior (equivalently, that $\{r_{n}>c(\alpha,y^{n},P)\}$ is dense in $\{r_{n}\geq c(\alpha,y^{n},P)\}$). the set convergence above is in the Hausdorff distance $d_{H}$; Otherwise a similar result can be proved with convergence taken in the Vietoris topology on closed subsets of $Y$.
Proof of Theorem 9.